OpenStudy (anonymous):

A round cookie cake, divided into 8 equal pieces, has a diameter of 16 inches. Marty eats two pieces of cake. What is the combined area (to the nearest tenth) of the two pieces that Marty eats? A)12.6 in2 B)27.8 in2 C)50.3 in2 D)201.1 in2

OpenStudy (anonymous):

I THINK B........... Anyway Welcome to OpenStudy, @taylormcclure !

pooja195 (pooja195):

C) Area of whole cake = pi * r^2, where r is the radius Area = pi * 8^2 = 64pi Marty eats 2/8 or 1/4 of the cake 64pi / 4 = 16pi = 50.3 square inches

OpenStudy (anonymous):

Oops, I was wrong...... :(

pooja195 (pooja195):

it ok

OpenStudy (anonymous):

thanks yall

pooja195 (pooja195):

np!!